Calculate the concentration of buffer components present in 263.00 mL of a buffer solution that contains 0.250 M NH4CL and 0.250 M NH2 immediately after the addition of 1.70 mL of 6.00 M HNO3 ?
Updated 42 mins ago:
What is
NH3 in M?
NH4 in M?
1
Expert's answer
2020-03-30T11:00:10-0400

n1 (NH3) = n1 (NH4+) = 263 × 0.250 / 1000 = 0.06575 mol

n (HNO3) = 6 × 1.7 / 1000 = 0.0102 mol

V2 = 263 + 1.7 = 264.7 mL = 0.2647 L

n2 (NH3) = n1 (NH3) - n (HNO3) = 0.06575 - 0.0102 = 0.05555 mol

n2 (NH4+) = n1 (NH4+) + n (HNO3) = 0.06575 + 0.0102 = 0.07595 mol

C = n / V

C (NH3) = 0.05555 / 0.2647 = 0.209 M

C (NH4+) = 0.07595 / 0.2647 = 0.287 M